is there a generation gap in the type of music we listen to?
Calculate the test statistic
find x^2stat & p value
using 0.05 level of significance
how do you use excel to solve?
media%2F190%2F1906bbca-d9a2-4b1d-bb3d-e7
Answer #1
Observed
72 67 40 36 215
43 19 8 2 72
50 46 48 32 176
43 52 62 75 232
19 29 36 42 126
17 26 38 48 129
6 11 18 15 50
250 250 250 250 1000
Expected
53.75 53.75 53.75 53.75 215
18 18 18 18 72
44 44 44 44 176
58 58 58 58 232
31.5 31.5 31.5 31.5 126
32.25 32.25 32.25 32.25 129
12.5 12.5 12.5 12.5 50
250 250 250 250 1000
p-value 3.20176E-17
TS 120.6226134

TS = 120.62

Formula used

Observed
72 67 40 36 =SUM(B2:E2)
43 19 8 2 =SUM(B3:E3)
50 46 48 32 =SUM(B4:E4)
43 52 62 75 =SUM(B5:E5)
19 29 36 42 =SUM(B6:E6)
17 26 38 48 =SUM(B7:E7)
6 11 18 15 =SUM(B8:E8)
=SUM(B2:B8) =SUM(C2:C8) =SUM(D2:D8) =SUM(E2:E8) =SUM(F2:F8)
Expected
=B$10*$F2/$F$10 =C$10*$F2/$F$10 =D$10*$F2/$F$10 =E$10*$F2/$F$10 =SUM(B12:E12)
=B$10*$F3/$F$10 =C$10*$F3/$F$10 =D$10*$F3/$F$10 =E$10*$F3/$F$10 =SUM(B13:E13)
=B$10*$F4/$F$10 =C$10*$F4/$F$10 =D$10*$F4/$F$10 =E$10*$F4/$F$10 =SUM(B14:E14)
=B$10*$F5/$F$10 =C$10*$F5/$F$10 =D$10*$F5/$F$10 =E$10*$F5/$F$10 =SUM(B15:E15)
=B$10*$F6/$F$10 =C$10*$F6/$F$10 =D$10*$F6/$F$10 =E$10*$F6/$F$10 =SUM(B16:E16)
=B$10*$F7/$F$10 =C$10*$F7/$F$10 =D$10*$F7/$F$10 =E$10*$F7/$F$10 =SUM(B17:E17)
=B$10*$F8/$F$10 =C$10*$F8/$F$10 =D$10*$F8/$F$10 =E$10*$F8/$F$10 =SUM(B18:E18)
=SUM(B12:B18) =SUM(C12:C18) =SUM(D12:D18) =SUM(E12:E18) =SUM(F12:F18)
p-value =CHISQ.TEST(B2:E8,B12:E18)
TS =CHISQ.INV.RT(B22,18)